OpenStudy (anonymous):

What is a polynomial function in standard form with zeroes 1, 2, 3, and –3?

OpenStudy (anonymous):

I have no idea how to do this

OpenStudy (retireed):

You have been given the roots. You just need to find "factors" that have cancelling values, when set equal to zero. Words I know that don't make much sense. I will simplify in the next example.

OpenStudy (retireed):

(x+3)(x-7) are factors of the polynomial x^2 - 4x -21 the ROOTs are x + 3 = 0 x = -3 x - 7 = 0 x = 7 the roots are -3, and 7 You were given the ROOTs you just need to work BACKWARDs ALSO you were given 4 ROOTs which will be more difficult when it comes to multiply the 3 or 4 factors out to get the polynomial in standard form.

OpenStudy (anonymous):

So would the answer be x^4-3x^3-7x^2+27x-18 = 0

OpenStudy (anonymous):

?

OpenStudy (retireed):

Why did I say 3 or 4 factors? Is because ONE of the factors is a double root. I think that is the term.

OpenStudy (retireed):

summer that could be the answer, but I haven't gotten that far yet. How did you get that answer? What factors did you multiply?

OpenStudy (anonymous):

3 and -3 is the double root then?

OpenStudy (retireed):

YES

OpenStudy (misty1212):

hmm

OpenStudy (retireed):

Factor 1 (x^2 - 3)

OpenStudy (misty1212):

not to butt in, but if 3 was a "double root' or a zeros of multiplicity two would have the factor \((x-3)^2\)

OpenStudy (anonymous):

So its not a double root?

OpenStudy (misty1212):

and if one factor was \(x^2-3\) then the zeros would be \(\sqrt3\) and \(-\sqrt3\)

OpenStudy (anonymous):

But was my answer right?

OpenStudy (misty1212):

idk i didn't do it, you want me to check?

OpenStudy (anonymous):

yes, I will also tell you how I did it

OpenStudy (misty1212):

your answer is correct

OpenStudy (retireed):

Misty you are right, I am working without paper, so my brain might not be seeing the whole picture. Thanks for your help.

OpenStudy (anonymous):

I tested multiplying by the factors which were 1,2,3,6,9,18. The coefficient is 1 and the trailing coefficient is -18.

OpenStudy (anonymous):

-3, 1,2, and 3 are the only ones who got zero as a result

OpenStudy (retireed):

Now I'm really confused, sorry.

OpenStudy (anonymous):

It's okay if you can't help, I can just ask my teacher

OpenStudy (anonymous):

Thanks anyways :)

OpenStudy (retireed):

Let's get the easy factors.

OpenStudy (retireed):

1, 2, 3, and –3 (x -1) (x -2) (x -3) (x +3) these makes more sense Can you do this? Multiply these out to get the standard form for the polynomial. My double root thought got me confused.